By the method of variation of parameters, solve the following differential equation:
d^2y/dx^2+4y =4 tan2x
1
Expert's answer
2020-11-22T18:26:34-0500

"\dfrac{d^2y}{dx^2} +4y = 4tan \ 2x\\
y''+4y = 4 tan \ 2x\\
y'' + 4y = 0\\
k^2+4 = 0\\
k_1 = 2i, \ k_2 = -2i\\
Therefore, \ an \ additional \ function \ is \ specified\\
y_c = C_1cos2x +C_2sin2x\\
y = Acos2x +Bsin2x\\
be \ the \ complete\ solution\\
 of\ the\ given\ equation\ where\\
 A\ and\ B\ are\ to\ be\ found\\
 We\ have\ y_1 = cos2x, \ y_2 = sin2x\\
y_1'=-2sin2x\\
y_2' = 2cos2x\\
Then\ W = y_1*y_2' -y_2*y_1' = \\
= 2cos^2\ 2x + 2sin^2\ 2x = 2\\
A' = \dfrac{-y_2 * 4tan\ 2x}{W}\\
B' = \dfrac{y_1 * 4tan\ 2x}{W}\\
A' = \dfrac{-sin\ 2x * 4tan\ 2x}{2}\\
B' = \dfrac{-cos\ 2x* 4tan\ 2x}{2}\\
A =\int \dfrac{-2sin^2\ 2x }{cos2x} dx\\ \ \\
B = \int 2sin\ 2x\ dx\\
A = -log(sec\ 2x +tan\ 2x) + sin\ 2x+C_1\\
B = -cos\ 2x + C_2\\
y = Acos2x +Bsin2x\\
y = C_1cos2x +C_2sin2x -cos\ 2x*log(sec\ 2x +tan\ 2x)\\

answer: \\
y = C_1cos2x +C_2sin2x -cos\ 2x*log(sec\ 2x +tan\ 2x)\\"